The vertex angle of an isosceles triangle is 72°,and each of equal sides is 10 cm.Find the perimeter and the area of of the triangle.
Answer by mananth(16949)

let the vertex be A where the legs meet
legs are 10 cm each
B & C are the vertices of the base
Law of Sines
a/sin A = b/sin B = c/sin C
a/sin 72 = 10/Sin 54 = c/Sin 54
a= 10*sin 72/Sin 54
a= 11.75
So the perimeter = 11.75+10+10
perimeter = 31.75 cm